# MVMoE: Multi-Task Vehicle Routing Solver with Mixture-of-Experts

### v1

@@ Abstract
Learning to solve vehicle routing problems (VRPs) has garnered much attention. However, most neural solvers are only structured and trained independently on a specific problem, making them less generic and practical. In this paper, we aim to develop a unified neural solver that can cope with a range of VRP variants simultaneously. Specifically, we propose a multi-task vehicle routing solver with mixture-of-experts (MVMoE), which greatly enhances the model capacity without a proportional increase in computation. We further develop a hierarchical gating mechanism for the MVMoE, delivering a good trade-off between empirical performance and computational complexity. Experimentally, our method significantly promotes the zero-shot generalization performance on 10 unseen VRP variants, and showcases decent results on the few-shot setting and real-world benchmark instances. We further provide extensive studies on the effect of MoE configurations in solving VRPs. Surprisingly, the hierarchical gating can achieve much better out-of-distribution generalization performance. The source code is available at: https://github.com/RoyalSkye/Routing-MVMoE.
@@ Acknowledgements
This work is supported by the funding (AISG3-RP-2022-031) awarded by AI Singapore, and the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant. We would like to thank the anonymous reviewers and (S)ACs of ICML 2024 for their constructive comments and dedicated service to the community. Jianan Zhou would like to personally express deep gratitude to his grandfather, Jinlong Hu, for his meticulous care and love during last 26 years. Eternal easy rest in sweet slumber.
@@ Appendix D Discussions
The motivation of multi-task VRP solver. 1) From the perspective of operation research (OR): We believe generality is always a favorable objective in the community. Both exact solvers (e.g., CPLEX, Gurobi) and heuristic solvers (e.g., LKH3) can solve a wide range of problems, making them a popular choice for solving VRPs (or COPs). Despite the recent trend of developing neural solvers, they still need to be tailored for each specific problem, which may consume much computation resources. For example, if training a simple model for each problem, the total training cost would be 16 models x 3 days per model, while the training cost of a unified model is only 3 days. This resource consumption cannot be neglected from a global and practical view, since we often have not much computational resources and time in practice to train many individual (and large already) models. 2) From the perspective of ML: The recent success of LLMs has demonstrated the power of DNNs. By only using text (and vision) prompts, they can solve a wide range of tasks, including code generation, text summary, or even solving optimization problems. It is also a research trend to develop a foundation model that can unify multiple tasks in CV and NLP communities. With that said, developing a more and more general neural solver (or foundation model) in combinatorial optimization is important, and we believe it should receive attention in the NCO community, and will be a trend in future research. Moreover, given a pretrained multi-task solver, we can also adapt it to the specific downstream task through efficient fine-tuning, making it an attractive choice compared to training from scratch.

### v2

@@ Abstract
Learning to solve vehicle routing problems (VRPs) has garnered much attention. However, most neural solvers are only structured and trained independently on a specific problem, making them less generic and practical. In this paper, we aim to develop a unified neural solver that can cope with a range of VRP variants simultaneously. Specifically, we propose a multi-task vehicle routing solver with mixture-of-experts (MVMoE), which greatly enhances the model capacity without a proportional increase in computation. We further develop a hierarchical gating mechanism for the MVMoE, delivering a good trade-off between empirical performance and computational complexity. Experimentally, our method significantly promotes zero-shot generalization performance on 10 unseen VRP variants, and showcases decent results on the few-shot setting and real-world benchmark instances. We further conduct extensive studies on the effect of MoE configurations in solving VRPs, and observe the superiority of hierarchical gating when facing out-of-distribution data. The source code is available at: https://github.com/RoyalSkye/Routing-MVMoE.
@@ Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-031), the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant, the National Natural Science Foundation of China (Grant 62102228), and the Natural Science Foundation of Shandong Province (Grant ZR2021QF063). We would like to thank the anonymous reviewers and (S)ACs of ICML 2024 for their constructive comments and dedicated service to the community. Jianan Zhou would like to personally express deep gratitude to his grandfather, Jinlong Hu, for his meticulous care and love during last 26 years. Eternal easy rest in sweet slumber.
@@ Appendix D Discussions
The motivation of multi-task VRP solver. 1) From the perspective of operation research (OR): We believe generality is a favorable objective in the community. Both exact solvers (e.g., CPLEX, Gurobi) and heuristic solvers (e.g., LKH3) can solve a wide range of problems, making them a popular choice for solving VRPs (or COPs). Despite the recent trend of developing neural solvers, they still need to be tailored for each specific problem, which may consume much computation resources. For example, if training a simple model for each problem, the total training cost would be 16 models x 3 days per model, while the training cost of a unified model is only 3 days. This resource consumption cannot be neglected from a global and practical view, since we often have not much computational resources and time in practice to train many individual (and large already) models. 2) From the perspective of ML: The recent success of LLMs has demonstrated the power of foundation models. By only using text (and vision) prompts, they can solve a wide range of tasks, including code generation, text summary, or even solving optimization problems. It is also a research trend to develop a foundation model that can unify multiple tasks in CV and NLP communities. With that said, developing a general neural solver (or foundation model) in combinatorial optimization is important, and we believe it should receive attention in the NCO community, and will be a trend in future research. Moreover, given a pretrained multi-task solver, we can also adapt it to the specific downstream task through in-context learning or efficient fine-tuning, making it an attractive choice versus training from scratch.